OpenStudy (anonymous):

A wheel has a radius of 1.4 feet and turns with an angular speed of 2.65 radians per second. Determine the distance traveled by a point on the wheel in 3 minutes.,

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

do u know ???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3 min = 3*60 = 180 s 2.65 radians per second = theta / time theta = 2.65* 180 s theta = 477 we can write 477 as 360 + 117

OpenStudy (anonymous):

convert theta into dist 360 =2pi and u will get ur answer

OpenStudy (anonymous):

360 = 2(pi)r sry typo

OpenStudy (anonymous):

ty i almost forgot that